OpenStudy (anonymous):

What is the sum of the first 35 consecutive odd numbers? Do i Multiply with this

OpenStudy (anonymous):

will it be 1225?

OpenStudy (zzr0ck3r):

what class is this for?

OpenStudy (zzr0ck3r):

\(\sum_1^{35}(2n-1)\)

OpenStudy (anonymous):

Math 104

OpenStudy (zzr0ck3r):

\(\sum_1^{35}(2n-1)=\sum_1^{35}2n-\sum_1^{35}1=2\sum_1^{35}n-35=2(630)-35=1225\)

OpenStudy (zzr0ck3r):

what is math 104?

OpenStudy (zzr0ck3r):

note: \(\sum_1^kn=\frac{k(k+1)}{2}\)

OpenStudy (anonymous):

ok thanks and I do this with the rest of my problems like this?

OpenStudy (zzr0ck3r):

no idea.

OpenStudy (zzr0ck3r):

all I really used was that last thing.

OpenStudy (zzr0ck3r):

\(\sum_1^kn=\frac{k(k+1)}{2}\)

OpenStudy (zzr0ck3r):

if you want to know why, look at example of summing up 1 - 20 take them and write them out like this 1 + 2 + 3 + 4 +5 + 6 + 7 + 8 + 9+10 +20+19+18+17+16+15+14+13+12+11 now if we add each column up and down we get 21+21+21+21+21+21+21+21+21+21 =10(21)=(20*21)/2

OpenStudy (jdoe0001):

tis just an arithmetic sequence all odd numbers up to 35 1, 3, 5, 7 .... then you find the "common difference" and then get the partial sum of them

OpenStudy (jdoe0001):

as zzr0ck3r showed above, using the \(\Large \Sigma\) notation

OpenStudy (anonymous):

ok thx